A company just paid an annual dividend of Rs.2. The dividend is expected to grow 30% per
year for the next two years and at a more sustainable growth rate thereafter. Two years later
the expected sales of the company are likely to be Rs.100 million and a net profit of Rs.10
million. The total Debt of the company is expected to be Rs.30 million and Total Equity of
Rs.10 million. The Company is likely to maintain a dividend pay-out ratio of 30% after 2
years and that is likely to continue forever. Find the current price of the stock which has a
beta of 1.2. The risk-free rate is 6% and the market risk premium is 6%.
1
Expert's answer
2020-12-17T04:44:41-0500

Solution to question #150254

Annual dividend of Rs.2. Growth is 30% per

year for the next two years and at a more sustainable growth rate thereafter. In two years:

The Company is likely to maintain a dividend payout ratio of 30% after 2 years and that is likely to continue forever. Find the current price of the stock which has a beta of 1.2. The risk-free rate is 6% and the market risk premium is 6%

"ER= RFR + 

Beta \times Market risk premium"


"Expected return= 6+1.2\times6= 13.2"


"Dividend year 1= D0\times(1+g) =Rs.2\times1.3 =Rs.2.6"

"Dividend year 2=D1\times(1+g) =2.6\times1.30 =Rs.3.38"

"Current price of stock=\frac{Dividend(1+g)}{r-g}"

where g is the growth rate and r is the expected return.

"Current price of stock=\frac{Rs.3.38}{0.132-0.3}"

Growth rate of 30% is greater than expected return 13.2%, this means that the discount model understates intrinsic value of the firm. The company is therefore a risk less arbitrage and will attract all money in the world.

N.B: The price cannot be computed because growth rate is higher than the expected return.